OpenStudy (jackellyn):

Write the following as a series that starts at n=0 instead of n=3.

OpenStudy (jackellyn):

\[\sum_{n=3}^{infinity}n ^{2}a _{n-1}(x+4)^{n+2}\]

OpenStudy (anonymous):

just put n=n+3

OpenStudy (anonymous):

clear right? where ever you see an \(n\) add 3 to it

OpenStudy (anonymous):

\[\sum_{0}^{infinity} (n+3)^{2} a _{n+2} (x+4)^{n+5}\]

OpenStudy (anonymous):

\infty

OpenStudy (anonymous):

\(\infty\)

OpenStudy (jackellyn):

Got it, thanks!